OpenStudy (anonymous):

4 · 9 − 4(3 · 4 − 11) ------------------ 4(3 + 2) − 3^2 + 5

OpenStudy (anonymous):

\[\frac{4 \times 9 - 4(3 \times 4-11)}{4(3+2)-3^{2}+5}\] \[\frac{36 - 4(12-11)}{4(5)-3^{2}+5}\] \[\frac{36 - 4}{20-9+5}\] \[\frac{32}{16} = 2\]

OpenStudy (anonymous):

Thank you!!!!! much appreciated

OpenStudy (anonymous):

A medal would be nice :)

OpenStudy (anonymous):

lol! If I had one I would deffinitlyy award you with it!

OpenStudy (anonymous):

Just click "Good Answer" next to the person's name to give medals

OpenStudy (anonymous):

i dont see where to click it :/

OpenStudy (anonymous):

my sreen doesnt have those buttons! thats dumb, its just blank white space

OpenStudy (anonymous):

ok! lol i just had to refresh my screen

OpenStudy (anonymous):

If you arent busy I got another one that would prolly be easy for you that i need answering... 3/8 × 16/3 + 1/2 × 14